OpenStudy (anonymous):

The table shows the height of a plant as it grows. What equation in point-slope form gives the plant’s height at any time? Let y stand for the height of the plant in cm and let x stand for the time in months. Time (months) Plant Height (cm) 3 15 5 25 7 35 9 45 A. y – 15 = (x – 3) B. y – 15 = 5(x – 3) C. y – 3 = (x – 15) D. The relationship cannot be modeled.

OpenStudy (imstuck):

The points are these: (3, 15), (5, 25), (7, 35), (9, 45), right?

OpenStudy (anonymous):

RITE

OpenStudy (imstuck):

In order to find out if these are linear, ie. they fit into a point-slope equation, they all have to fall onto a straight line. In order to do that, we should probably graph it. Hold on for sec, ok?

OpenStudy (anonymous):

KKK

OpenStudy (imstuck):

I'm trying to do something here...

OpenStudy (imstuck):

Here is the pic of three of those points on a piece of graph paper. As you can see, they do indeed fall in a straight line, right? So we can, in fact, find an equation for it!

OpenStudy (imstuck):

Take 2 of the points and first find the slope of the line using the formula:\[\frac{y _{2} -y _{1}}{x _{2}-x _{1} }\]

OpenStudy (imstuck):

Filling in using the first two points in your table, (3, 15) and (5, 25) we get:

OpenStudy (anonymous):

hey my quiz is on a timer i have 2 min left so plz

OpenStudy (anonymous):

u no what im done u took to long

OpenStudy (imstuck):

\[\frac{ 25-15 }{ 5-3 }=\frac{ 10 }{ 2 }=5\]So the slope of the line is 5. The standard form for an equation in point-slope form is\[y - y _{1}= m(x-x _{1})\]where m is the slope you just found, and the x1 and y1 are points from one of your given coordinates. Again, let us use the first point of (3, 15) for x1 and y1 to get this:

OpenStudy (imstuck):

sorry...actually we are not allowed to help on tests or quizzes here!

OpenStudy (imstuck):

But the answer, just because is y = 5x